20160315MEMO » History » Revision 1

Revision 1/2 | Next »
Boyd, Steven, 15 March 2016 13:22

Operations report to MEMO

  • Data-taking
    • LiH absorber was installed last week with no problems and we are now taking data smoothly
    • pressure in decay solenoid fridge is still oscillating. Mike/Mark are aware of it.
    • we've negotiated an increase in ISIS beamloss trip levels to 2.5 This should allow
      us to run at 2.0 rather than 1.3 (which is what we've been averaging) and help with the data rate.
    • Will continue taking LiH data to the end of the user cycle.
  • User cycle 2016/01
    • Next user cycle starts on April 12th
    • We've decided that this will mostly be taken up by Hall activities (QP/QD work, realignment work, etc)
    • We do want to bolster the zero absorber data-set though, and will use two (possibly three) weekends in
      April to do so.
    • I'm putting staffing in place for this - and will send out a shift email soon.
  • Cycle 2016/02 : (June/July)
    • What we do here very much depends on what is ready
    • The current plan is to take energy loss data with LiH (or vacuum) and magnetic fields (SSU (E-C-E) + FC + SSD (E-C-E))
    • We need to understand if (and when) we power M2 in SSD for a cooling measurement

Updated by Boyd, Steven over 8 years ago · 1 revisions